OpenStudy (anonymous):

how does \[r^2+2 r+\lambda = 0\] come out to be \[-1\pm \sqrt{1-\lambda}\] I got \[-1\pm \sqrt{4-4\lambda} = 1\pm2\sqrt{1-\lambda}\]

OpenStudy (anonymous):

*edit- the last equation should be -1 plus/minus ...

OpenStudy (anonymous):

\[r^2 + 2r + 1 - 1 + \lambda= 0 \] \[(r + 1)^2 = 1 - \lambda\] \[r + 1 = \pm\sqrt{1 - \lambda}\] \[r = -1 \pm \sqrt{1 - \lambda}\]

OpenStudy (anonymous):

wow, forgot to divide by 2 for both real AND imaginary parts using quadratic eq.. thanks!

OpenStudy (anonymous):

No Problem :)